OpenStudy (anonymous):

will fan and medal !!! Alfred is taking a trip. He drove 150 mi and then stopped to refuel. He will drive 60 mph to finish his trip. Let x represent the hours it takes for Alfred to complete the rest of his trip, and let y represent the total distance traveled. Complete the function equation that relates Alfred’s total distance traveled to his time.

Directrix (directrix):

For the rest of Alfred's trip after the refuel stop, the distance will be his rate of speed times the number of hours x it takes for him to complete the trip. If y represents the total distance traveled, then: y = 150 + 60*x Agree? @OGangsta

OpenStudy (anonymous):

it is supposed to be on point slope form though @KamiBug pls help

OpenStudy (kamibug):

y = 60x + 150 Just rearrange it. :)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

can you help me with one more please

OpenStudy (kamibug):

Sure :)

OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the point (−1, 2) and has a slope of 0.9.

OpenStudy (kamibug):

Plug the given info into the point-slope form. Then, solve for y to get it in slope-intercept form. :) y-2=0.9(x+1) <--- point-slope form Distribute. y-2=0.9x + 0.9 Combine Like Terms. y=0.9x + 2.9 <--- slope-intercept form

OpenStudy (anonymous):

thank you so much no wonder why you were my first fan

OpenStudy (kamibug):

Lol, no problem. :)